OpenStudy (anonymous):

Aimi is making Valentine’s Day cards for everyone in her class. She plans to use a whole sheet of paper for each of her 6 closest friends. She will use 1/8 of a sheet of paper for everyone else in the class. She has 34 sheets of paper. How many of the smaller cards can she make? A. 3 cards B. 4 cards C. 272 cards D. 224 cards

OpenStudy (australopithecus):

6 sheets of paper (1/8) sheet of paper for every one in class 34 sheets of paper Total amount of paper - Sheets of Paper for Best Friend Cards = Amount of Paper left over so, 34 - 6 = 28 so she has 28 sheets of paper after she makes cards for her friends she is using 1/8 of a sheet so per piece of paper so so, (1/8)x = 18 sheet where x = the amount of cards she can make

OpenStudy (australopithecus):

that last part should be, so, (1/8)x = 28 sheets where x = the amount of cards she can make

OpenStudy (anonymous):

thant makes no sense

OpenStudy (australopithecus):

what doesn't make sense about it?

OpenStudy (anonymous):

honestly the whole thing ive never done any of this its my first time

OpenStudy (australopithecus):

well think about what multiplication really is. (1/8)x = 1 sheet x = 8 cards

OpenStudy (australopithecus):

and go over it until it makes sense to you.

OpenStudy (anonymous):

yea thats not going to help going over it i have a problem understanding math i can hardly do my multiplcation

OpenStudy (anonymous):

@Australopithecus

OpenStudy (australopithecus):

well I recommend doing some reading up on it, there are others out there who can explain math better than I and more correctly. Spend the time learning it, what one man can devise another can learn with diligence and discipline.